November 2, 2017: Lund update

First info (VREF, noise, gain) on a few tested V3 and V4 chips circulated to the email-list

by Anders (Oct 27 & Oct 31). So far, noise and gain etc look quite similar to V2 chips. More statistics and detailed info will come shortly/ASAP.

Plots collected under: http://www.hep.lu.se/staff/silvermyr/SAMPA/calib/

In these slides : A first look at odd-even effects for V3 chips

Anders, David, Jonatan, Lennart, Ulf

As before, we insert a rather big pulse (close to 1000 ADC counts), and then look at the distribution of ADC values for all samples in many events.

ADC distributions (ADC range 900 - 910), few channels 3

Next, we zoom in to a range of higher ADC values: 900-910 (range chosen somewhat arbitrarily).

ADC distributions (ADC range 900 - 910), all channels 4

We can also look at the data for all 32 channels - more such plots are available in the backup.

ADC distributions (ADC range 500 - 510), few channels 5

Next, we zoom in to a range of middle ADC values: 500-510 (also chosen somewhat arbitrarily). Somewhat of an odd/even effect also in that range?

At this point, seeing what looked like some odd/even pattern, at least at higher ADC values, we worried about if our way of sending the pulses (at regular time intervals) could maybe help create an odd/even pattern?

To investigate this, we made a more random pulse generator yesterday (based on a radioactive source and a PM tube/counter), and that is what is used for the following slides...

As before, we insert a rather big pulse (close to 1000 ADC counts), and then look at the distribution of ADC values for all samples in many events.

Random pulses, (ADC range 900 - 910), few channels 8

Next, we zoom in to a range of higher ADC values: 900-910 (range chosen somewhat arbitrarily).

Random pulses, (ADC range 900 - 910), all channels 9

We can also look at the data for all 32 channels - more such plots are available in the backup.

Next, we zoom in to a range of middle ADC values: 500-510 (chosen somewhat arbitrarily). Somewhat of an odd/even effect in that range too?

Summary / Outlook 11

Odd/even effect still present for V3. Will check V4 next.

Histogram and text dump info available (for Armando et al) under: http://www.hep.lu.se/staff/silvermyr/SAMPA/odd even/V3/ e.g. http://www.hep.lu.se/staff/silvermyr/SAMPA/odd even/V3/regularpulse/dump.txt http://www.hep.lu.se/staff/silvermyr/SAMPA/odd even/V3/regularpulse/dump verbose.txt http://www.hep.lu.se/staff/silvermyr/SAMPA/odd even/V3/randompulse/dump.txt http://www.hep.lu.se/staff/silvermyr/SAMPA/odd even/V3/randompulse/dump verbose.txt

Will continue to look at more chips...

Backup 12

More plots with all 32 channels included...

ADC distributions (ADC range 100 - 1000), all channels 13 ADC distributions (ADC range 100 - 110), all channels 14 ADC distributions (ADC range 500 - 510), all channels 15 ADC distributions (ADC range 900 - 910), all channels 16 Random pulses, (ADC range 100 - 1000), all channels 17 Random pulses, (ADC range 100 - 110), all channels 18 Random pulses, (ADC range 500 - 510), all channels 19 Random pulses, (ADC range 900 - 910), all channels 20